The area of a rectangle is 20. The length is 3 less than twice the width. Find the length and width.
Answer by vksarvepalli(154)

let the width of the rectangle be w
The length is 3 less than twice the width
so the length = 2w-3
given area = 20
so w(2w-3)=20
so 2w^2-3w-20=0
=> 2w^2-8w+5w-20=0
=> 2w(w-4)+5(w-4)=0
so (2w+5)(w-4)=0
so w=4 or w=-5/2
since width can't be negative
the width of the rectangle is = 4
so the length will be = 5
Ans: length=5 width=4